# Identifying non-Hermitian critical points with quantum metric

### v1

@@ III.2 Revealing mobility edges
Here $g$ represents the nonreciprocal strength of the hopping amplitude, $\Delta$ is the amplitude of irrational modulation with $\beta=(\sqrt{5}-1)/2$, and $\alpha\in(-1,1)$ denotes the lattice parameter. In the Hermitian case ($g=0$), when $\alpha=0$, it reduces to the standard AA model without mobility edges. When $\alpha\neq 0$, this GAA model in the Hermitian limit has an exact mobility edge at energy $E=E_{c}$ that follows the relationship Ganeshan et al. (2015)
@@
The presence of the mobility edge can be understood by the energy-independent localization transition and self-duality condition driven by the cosine dispersion Ganeshan et al. (2015). This GAA model has been realized with ultracold atoms in tunable synthetic lattices and the mobility edge has been observed An et al. (2021); Wang et al. (2022). However, the mobility edge in the non-Hermitian GAA model is yet to be explored.
To characterize the localization properties of all eigenstates and reveal the mobility edge in the Hermitian and non-Hermitian cases, we can numerically compute the participation ratio of the $n$-th eigenstate $|\psi_{n}\rangle$:
$\mathrm{PR}_{n}=\frac{1}{L}\frac{1}{\sum_{j}^{L}|\braket{j}{\psi_{n}}|^{2}}.$ (13)
The extended and localized eigenstates take $\mathrm{PR}_{n}\to 1$ and $\mathrm{PR}_{n}\to 0$, respectively. To detect the mobility edge with the quantum metric, we numerically calculate the corresponding diagonal element of the quantum metric $g^{(n)}_{\mu\mu}$ with $\mu=\{\Delta,\alpha\}$ for the $n$-th eigenstate $|\psi_{n}\rangle$.
We first show the numerical results in the Hermitian limit with $g=0$ in Fig. 3. As shown in Fig. 3(a), the mobility edge exhibits and separates the eigenstates from $\mathrm{PR}_{n}$: the upper ones as extended states while the lower ones are localized states. The peak positions of $g^{(n)}_{\Delta\Delta}$ and the exact mobility edge given by Eq. (12) are plotted, which are consistent with each other. To be more clearly, we show $g^{(n)}_{\Delta\Delta}$ and $\mathrm{PR}_{n}$ as a function of $\Delta$ for the ground state ($n=1$) and the highest excited state ($n=L$) in Fig. 3(b), where $g^{(n)}_{\Delta\Delta}$ exhibits a distinct peak at the critical point with $\mathrm{PR}_{n}\approx 0$. The mobility edge separating extended and localized eigenstates with respect to varying the parameter $\alpha$ can be also revealed by the quantum metric $g^{(n)}_{\alpha\alpha}$, as shown in Figs. 3(c,d).
@@ IV.1 Non-Hermitian cluster Ising model
We proceed to reveal the quantum phase transition of many-body ground states in non-Hermitian interacting spin systems with the quantum metric. We first consider the non-Hermitian cluster Ising model with exactly solvable ground states Guo et al. (2022), which takes the following Hamiltonian
@@
Using the standard Jordan-Wigner transformation and the Fourier transformation under the PBC, we can obtain the Hamiltonian in momentum space, which is given by
$H_{k}=2\sum_{k>0}[iy_{k}(c_{k}^{\dagger}c_{-k}^{\dagger}+c_{k}c_{-k})+z_{k}(c_{k}^{\dagger}c_{k}+c_{-k}^{\dagger}c_{-k}-1)],$ (15)
here $y_{k}=\sin(2k)+\lambda\sin(k)$ and $z_{k}=\cos(2k)-\lambda\cos(k)-\frac{i\Gamma}{4}$. By diagonalizing $H_{k}$ through the Bogoliubov transformation, one can obtain the ground state of this model
|$\ket{G}=\prod_{k>0}[\cos(\frac{\theta_{k}}{2})+i\sin(\frac{\theta_{k}}{2})c_{k}^{\dagger}c_{-k}^{\dagger}]\ket{\mathrm{Vac}},$ (16)
where |$\ket{\mathrm{Vac}}$ is the vacuum state of the free fermion and $\theta_{k}=\arctan(-y_{k}/z_{k})$. The properties of the ground state can be characterized by the staggered magnetization $m_{y}$ and the string order parameter $O^{x}$. The staggered magnetization $m_{y}$ is determined by the spin-correlation function $R_{r}^{y}$ with the following relationship
$m_{y}^{2}=\lim_{r\to\infty}(-1)^{r}R_{r}^{y}=\lim_{r\to\infty}(-1)^{r}\braket{G}{\sigma_{l+r}^{y}\sigma_{l}^{y}}{G},$ (17)
where $r$ is the relative distance between spins and $l$ ($l\in[1,N]$) is an irrelevant index under the PBC. Using the Wick theorem Wick (1950), $R_{r}^{y}$ can be transformed to a Toeplitz determinant
$R^{y}_{r}=\begin{vmatrix}D_{1}&D_{0}&\dots&D_{-r+2}\\ D_{2}&D_{1}&\dots&D_{-r+3}\\ \vdots&\vdots&\ddots&\vdots\\ D_{r}&D_{r-1}&\dots&D_{1}\end{vmatrix},$ (18)
with the elements
@@
The second order parameter is the string order parameter $O^{x}$, which is given by
$O^{x}=\lim_{r\to\infty}(-1)^{r}\langle G|\sigma_{1}^{x}\sigma_{2}^{y}(\prod_{l=3}^{r}\sigma^{z}_{l})\sigma_{r+1}^{y}\sigma_{r+2}^{x}|G\rangle.$ (20)
Using the same method in Ref. Wick (1950), one can rewrite $O^{x}$ as
$O^{x}=\lim_{r\to\infty}\begin{vmatrix}G_{-2}&G_{-3}&\dots&G_{-r-1}\\ G_{-1}&G_{-2}&\dots&G_{-r}\\ G_{0}&G_{-1}&\dots&G_{-r+1}\\ \vdots&\vdots&\ddots&\vdots\\ G_{r-2}&G_{r-5}&\dots&G_{-3}\\ G_{r-3}&G_{r-4}&\dots&G_{-2}\end{vmatrix},$ (21)
@@
Figure 5: (Color online) The non-Hermitian cluster Ising model $H_{\text{NH-Ising}}^{(1)}$ under the PBC. (a) The phase diagram with boundaries (blue solid lines) determined by the order parameters $O^{x}$ and $m_{y}$. Four regions $\mathrm{I}$, $\mathrm{II}$, $\mathrm{III}$, $\mathrm{IV}$ denote the cluster, gapless, paramagnetic, and antiferromagnetic phases, respectively. The locations of the peaks of the quantum metric are plotted as the red circles and consistent with the phase boundaries. (b) $g_{\Gamma\Gamma}$ as a function of $\Gamma$ with fixed $\lambda=0.5$ and $\lambda=0.9$. (c) In the Hermitian limit with $\Gamma=0$ and (d) in the non-Hermitian case with $\Gamma=2$, $g_{\lambda\lambda}$, $O^{x}$, and $m_{y}$ as a function of $\lambda$. Other parameter is $J=1$.
By numerically calculating the order parameters $O^{x}$ and $m_{y}$ for $r=1000$ that is large enough to approach the thermodynamic limit, one can obtain four phases in the phase diagram in the $\lambda$-$\Gamma$ plane Guo et al. (2022), as shown in Fig. 5(a). The first one is the gapped cluster phase with long-range correlation ($O^{x}\neq 0$) and vanishing staggered magnetization ($m_{y}=0$). The second one is the gapless phase with the power-law decay of correlation ($O^{x}\neq 0$ for finite systems) and vanishing staggered magnetization ($m_{y}=0$). The third and fourth ones are the conventional paramagnetic ($m_{y}=0$) and antiferromagnetic ($m_{y}\neq 0$) without long-range correlation, respectively. The phase boundaries determined by the order parameters are plotted as the blue solid lines in Fig. 5(a). To show that the quantum phase transition points in this non-Hermitian cluster Ising model can be revealed by the quantum metric, we plot the locations of the peaks of $g_{\mu\mu}$ of the ground state |$\ket{G}$ for the parameter $\mu=\{\lambda,\Gamma\}$. As shown in Fig. 5(a), the locations of the peaks correspond to the critical points of the four different phases. To illustrate more clearly, we plot $g_{\Gamma\Gamma}$ as a function of $\Gamma$ in Fig. 5(b). One can still find the peaks at $\Gamma\approx 2.23$ with fixed $\lambda=0.5$, and at $\Gamma\approx 0.48$ and $\Gamma\approx 2.95$ with fixed $\lambda=0.9$ in $g_{\Gamma\Gamma}$, indicating one and two quantum critical points, respectively. We further plot $g_{\lambda\lambda}$ as a function of $\lambda$ in the Hermitian ($\Gamma=0$) and non-Hermitian case ($\Gamma=2$) in Fig. 5(c) and Fig. 5(d), respectively. The results show that the values of $g_{\lambda\lambda}$ exhibit peaks at the quantum phase transition points where the behaviours of the order parameters $O^{x}$ and $m_{y}$ change.
@@
Figure 6: (Color online) The non-Hermitian mixed-field Ising model $H_{\text{NH-Ising}}^{(2)}$ with the system size $N=10$. (a) $|M_{z}|$ and (b) $\xi_{h_{z}}$ of the ground state as functions of $h_{z}$ and $h_{x}$ under periodic boundary condition. The phase diagram based on $|M_{z}|$ consists two regions $\mathrm{I}$ and $\mathrm{II}$ for the paramagnetic and ferromagnetic phases with real and complex energies, respectively. $|M_{z}|$ and $\xi_{h_{z}}$ as a function of $h_{z}$ (c) under the PBC; and (d) the OBC, respectively. Other parameter is $J=1$.

### v3

@@ III.2 Revealing mobility edges
Here $g$ represents the nonreciprocal strength of the hopping amplitude, $\Delta$ is the amplitude of irrational modulation with $\beta=(\sqrt{5}-1)/2$, and $\alpha\in(-1,1)$ denotes the lattice parameter. Here we also choose the lattice size $L=610$ from the Fibonacci sequence to satisfy the PBC. It is sufficiently large for $H_{\text{GAA}}^{(2)}$ and the finite-size scaling will be given in Sec. V.
In the Hermitian case ($g=0$), when $\alpha=0$, it reduces to the standard AA model without mobility edges. When $\alpha\neq 0$, this GAA model in the Hermitian limit has an exact mobility edge at energy $E=E_{c}$ that follows the relationship Ganeshan et al. (2015)
@@
The presence of the mobility edge can be understood by the energy-independent localization transition and self-duality condition driven by the cosine dispersion Ganeshan et al. (2015); Li and Li (2024). This GAA model has been realized with ultracold atoms in tunable synthetic lattices and the mobility edge has been observed An et al. (2021); Wang et al. (2022). However, the mobility edge in the non-Hermitian GAA model is yet to be explored. To characterize the localization properties of all eigenstates and reveal the mobility edge in the Hermitian and non-Hermitian cases, we can numerically compute the participation ratio of the $n$-th eigenstate $|\psi_{n}\rangle$:
$\mathrm{PR}_{n}=\frac{1}{L}\frac{1}{\sum_{j}^{L}|\braket{j}{\psi_{n}}|^{4}}.$ (13)
The extended and localized eigenstates take $\mathrm{PR}_{n}\to 1$ and $\mathrm{PR}_{n}\to 0$, respectively.
To detect the mobility edge with the quantum metric, we numerically calculate the diagonal elements of the quantum metric $g^{(n)}_{\mu\mu}$ with $\mu=\{\Delta,\alpha\}$ for the $n$-th eigenstate $|\psi_{n}\rangle$, which are equivalent to the fidelity susceptibility of $|\psi_{n}\rangle$ with respect to the parameter $\mu$. We first show the numerical results in the Hermitian limit with $g=0$ in Fig. 3. As shown in Fig. 3(a), the mobility edge exhibits and separates the eigenstates from $\mathrm{PR}_{n}$: the upper ones as extended states while the lower ones are localized states. The peak positions of $g^{(n)}_{\Delta\Delta}$ and the exact mobility edge given by Eq. (12) are plotted, which are consistent with each other. To be more clearly, we show $g^{(n)}_{\Delta\Delta}$ and $\mathrm{PR}_{n}$ as a function of $\Delta$ for the ground state ($n=1$) and the highest excited state ($n=L$) in Fig. 3(b), where $g^{(n)}_{\Delta\Delta}$ exhibits a distinct peak at the critical point with $\mathrm{PR}_{n}\approx 0$. The mobility edge separating extended and localized eigenstates with respect to varying the parameter $\alpha$ can be also revealed by the quantum metric $g^{(n)}_{\alpha\alpha}$, as shown in Figs. 3(c,d).
@@ IV Quantum phase transitions in non-hermitian many-body systems
Figure 5: (Color online)The non-Hermitian cluster Ising model $H_{\text{NH-Ising}}^{(1)}$ under the PBC. (a)$\Delta_{R}$ and (b) $\Delta_{I}$ as functions of $\lambda$ and $\Gamma$, respectively. The parameters ($\lambda$, $\Gamma$) for the markers in (a-f) are (0.1,3) for the blue diamond, (0.5,3) for the green square, (0.96,3) for the pink up-triangle, (1.4,3) for the purple pentagram, and (1.9,3) for the orange inverted triangle, respectively. (c) $|\Delta E_{R}|$ and $|\Delta E_{I}|$ versus $k$ for $\lambda=0.5$ (upper) and $\lambda=0.96$ (lower) with fixed $\Gamma=3$. (d) $|O_{r}^{x}|$ and (e) $|R_{r}^{y}|$ as a function of $r$, respectively. (f) $\mathrm{ln}|O_{r}^{x}|$ and $\mathrm{ln}|R_{r}^{y}|$ versus $\mathrm{ln}r$ with the power-law decay behaviour. (g) The phase diagram with boundaries determined from the order parameters (blue solid points) and the quantum metric (red circles) of the ground state, respectively. The regions from I to V correspond to five phases: The real gapped cluster phase in region I, the gapless quasi cluster phase in region II, the imaginary gapped quasi cluster phase in region III, the imaginary gapped antiferromagnetic phase in region $\mathrm{IV}$, and the real gapped antiferromagnetic phase in region $\mathrm{V}$, respectively. $g_{\lambda\lambda}$, $\partial O_{x}/\partial\lambda$, and $\partial m_{y}/\partial\lambda$ as a function of $\lambda$ for (h) $\Gamma=0$ and (i) $\Gamma=3$, respectively. (j) $g_{\Gamma\Gamma}$ as a function of $\Gamma$ with $\lambda=0.5,0.9$. Other parameter is $J=1$.
In this section, we proceed to reveal the quantum phase transition of many-body ground states in non-Hermitian interacting spin systems with the diagonal elements of quantum metric, which correspond to the ground-state fidelity susceptibility. Although most of experimental studies of non-Hermitian physics focus on single-particle systems, the extension to non-Hermitian many-body systems has attracted increasing attentions. For instance, the experimental scheme to realize the non-Hermitian mixed-filed Ising model by using Floquet Rydberg atomic arrays has been proposed  Shen et al. (2023). The cluster-Ising interaction and controllable dispassion can be experimentally achieved in Floquet superconducting qubit arrays Zhang et al. (2022).
@@ IV.1 Non-Hermitian cluster Ising model
We first consider the non-Hermitian cluster Ising model with exactly solvable ground states Guo et al. (2022), which takes the following Hamiltonian
@@
Using the standard Jordan-Wigner transformation and the Fourier transformation under the PBC, we can obtain the Hamiltonian in momentum space
$H_{k}=2\sum_{k=0}^{\pi}[iy_{k}(c_{k}^{\dagger}c_{-k}^{\dagger}+c_{k}c_{-k})+z_{k}(c_{k}^{\dagger}c_{k}+c_{-k}^{\dagger}c_{-k}-1)],$ (15)
where $y_{k}=\sin(2k)+\lambda\sin(k)$ is real and $z_{k}=\cos(2k)-\lambda\cos(k)-\frac{i\Gamma}{4}$ is generally complex. The Hamiltonian in Eq. (15) can be rewritten in the terms of the Bogoliubov-de-Gennes formalism
$H_{k}=\sum_{k=0}^{\pi}\begin{pmatrix}c_{k}^{\dagger}&c_{-k}\end{pmatrix}\begin{pmatrix}z_{k}&iy_{k}\\ -iy_{k}&-z_{k}\end{pmatrix}\begin{pmatrix}c_{k}\\ c_{-k}^{\dagger}\end{pmatrix}.$ (16)
The eigen-energies are given by $E_{k,\pm}=\pm\sqrt{z_{k}^{2}+y_{k}^{2}}$ as a function of momentum $k$, where $E_{k,-}$ and $E_{k,+}$ correspond to the ground-state and excitation spectra, respectively. In order to diagonalize $H_{k}$, one can take the non-Hermitian Bogoliubov transformation Lee and Chan (2014),
$\begin{split}\eta_{k}=u_{k}c_{k}+v_{k}c_{-k}^{\dagger},~{}~{}\eta_{-k}=-v_{k}c_{k}^{\dagger}+u_{k}c_{-k},\\ \bar{\eta}_{k}=u_{k}c_{k}^{\dagger}+v_{k}c_{-k},~{}~{}\bar{\eta}_{-k}=-v_{k}c_{k}+u_{k}c_{-k}^{\dagger},\end{split}$ (17)
where the complex factors $u_{k}$ and $v_{k}$ are given by
$\begin{split}u_{k}=\frac{z_{k}+E_{k,-}}{\mathcal{C}},~{}~{}v_{k}=-\frac{y_{k}}{\mathcal{C}},\end{split}$ (18)
with $\mathcal{C}$ being the constant to satisfy $u_{k}^{2}+v_{k}^{2}=1$. By diagonalizing $H_{k}$, we obtain the ground state of $H_{\text{NH-Ising}}^{(1)}$ as
|$\ket{G}=\frac{1}{\mathcal{N}}\prod_{k=0}^{\pi}[u_{k}-v_{k}c_{k}^{\dagger}c_{-k}^{\dagger}]\ket{\mathrm{0}},$ (19)
where $\mathcal{N}=\prod_{k=0}^{\pi}(|u_{k}|^{2}+|v_{k}|^{2})$ is the normalization constant, and |$\ket{\mathrm{0}}$ denotes the vacuum state without fermion occupied and is defined via $c_{k}\ket{\mathrm{0}}=0$.
The properties of the ground state can be characterized by the staggered magnetization $m_{y}$ and the string order parameter $O_{x}$ Smacchia et al. (2011); Guo et al. (2022). The staggered magnetization is $m_{y}^{2}=\lim_{r\to\infty}(-1)^{r}R_{r}^{y}$, where the two-spin correlation $R_{r}^{y}$ is given by Smacchia et al. (2011)
$\begin{split}R_{r}^{y}=\braket{\sigma_{l}^{y}\sigma_{l+r}^{y}}=(-1)^{r}\braket{A_{l}B_{l+1}A_{l+1}\dots A_{l+r-1}B_{l+r}}.\end{split}$ (20)
Here $A_{l}=c_{l}^{\dagger}+c_{l}$, $B_{l}=c_{l}^{\dagger}-c_{l}$, $r$ is the relative distance between two spins, and $l$ ($l\in[1,N]$) is an irrelevant index under the PBC. Using the Wick theorem Wick (1950), $R_{r}^{y}$ can be transformed to the Pfaffian of a $2r\times 2r$ skew-symmetric matrix Lee and Chan (2014)
$R^{y}_{r}=(-1)^{r}\rm{pf}\begin{vmatrix}0&-G_{-1}&Q_{1}&-G_{-2}&Q_{2}&\dots&-G_{-r}\\ &0&G_{0}&S_{1}&G_{1}&\dots&S_{r-1}\\ &&0&-G_{-1}&Q_{1}&\dots&-G_{-r+1}\\ &&&\ddots&\ddots&&\vdots&\\ &&&&0&G_{0}&S_{1}\\ &&&&&0&-G_{-1}\\ &&&&&&0\end{vmatrix},$ (21)
where $\rm{pf}(D)^{2}=\rm{det}(D)$ and its positive value is chosen in our calculation. The matrix elements $G_{r}=\braket{B_{l}A_{l+r}}=-\braket{A_{l+r}B_{l}}$, $S_{r}=\braket{B_{l}B_{l+r}}$, and $Q_{r}=\braket{A_{l}A_{l+r}}$ can be derived based on $u_{k}$ and $v_{k}$. By taking similar calculations as those given the appendices in Refs. Lee and Chan (2014); Zhao et al. (2022), one obtains the expressions of the matrix elements
@@
and
@@
respectively. The second order parameter is the string order $O_{x}=\lim_{r\to\infty}(-1)^{r}O^{x}_{r}$, where the string correlation is given by Smacchia et al. (2011)
$\begin{split}O^{x}_{r}&=\langle\sigma_{1}^{x}\sigma_{2}^{y}(\prod_{l=3}^{r}\sigma^{z}_{l})\sigma_{r+1}^{y}\sigma_{r+2}^{x}\rangle\\ &=\braket{B_{1}B_{2}A_{3}B_{3}\dots A_{r}B_{r}A_{r+1}A_{r+2}}.\end{split}$ (24)
Using the same method in Ref. Wick (1950), one can rewrite $O_{r}^{x}$ as
$O^{x}_{r}=\rm{pf}\begin{vmatrix}0&S_{1}&G_{2}&S_{2}&G_{3}&\dots&G_{r+1}\\ &0&G_{1}&S_{1}&G_{2}&\dots&G_{r}\\ &&0&-G_{0}&Q_{1}&\dots&Q_{r-1}\\ &&&\ddots&\ddots&&\vdots&\\ &&&&0&G_{1}&G_{2}\\ &&&&&0&Q_{1}\\ &&&&&&0\end{vmatrix}.$ (25)
We first study the energy gap of the ground state by using the complex energy spectra $\Delta E(k)=E_{k,+}-E_{k,-}\equiv\Delta E_{R}(k)+i\Delta E_{I}(k)$. The corresponding real and imaginary parts of the complex energy gap are given by $\Delta_{R}=\min_{k\in[0,\pi]}|\Delta E_{R}(k)|$ and $\Delta_{I}=\min_{k\in[0,\pi]}|\Delta E_{I}(k)|$, respectively. We numerically calculate $\Delta_{R}$ and $\Delta_{I}$ in the $\lambda$-$\Gamma$ plane, as shown in Figs. 5(a) and 5(b), respectively. One can find a parameter regime for the gapless phase with $\Delta_{R}=\Delta_{I}=0$, while other regimes for gapped phases with $\Delta_{R}\neq 0$ (real gap) or $\Delta_{I}\neq 0$ (imaginary gap) , corresponding to line gaps in the complex energy spectra. To be more clearly, we show two typical results of $|\Delta E_{R}|$ and $|\Delta E_{I}|$ as functions of $k$ in Fig. 5(c), where the upper and lower sub-figures indicate the gapless and imaginary gapped phases, respectively. We then study the order parameters of the ground state. Typical numerical results of $|O_{r}^{x}|$ and $|R_{r}^{y}|$ as functions of the distance $r$ are shown in Figs. 5(d) and 5(e), respectively. Several cases for the power-law decay of $|O_{r}^{x}|$ and $|R_{r}^{y}|$ with respect to $r$ are further shown in Fig. 5(f).
Combing the properties of energy gaps and order parameters, one can obtain the ground-state phase diagram of the non-Hermitian cluster Ising model Guo et al. (2022), as shown in Fig. 5(g). The region $\mathrm{I}$ corresponds the real gapped cluster phase with $\Delta_{R}\neq 0$ and long-range string correlation $|O_{r}^{x}|$ and $O_{x}\neq 0$ in the large $r$ limit. In regions $\mathrm{II}$ and $\mathrm{III}$, the ground states are respectively in the gapless and imaginary gapped cluster phases with quasi long-range string correlation, where $|O_{r}^{x}|$ exhibits the power-law decay as shown in Fig. 5(f). Note that the region III is overlooked in Ref. Guo et al. (2022) due to some errors in the BdG transformations and the calculation of long-range correlations. The phase in the region $\mathrm{IV}$ is the imaginary gapped antiferromagnetic phase with quasi long-range correlation of $R_{r}^{y}$ ($|R_{r}^{y}|$ exhibits the power-law decay). The region $\mathrm{V}$ corresponds to the real gapped antiferromagnetic phase with $\Delta_{R}\neq 0$ and long-range correlation of $|R_{r}^{y}|$ and $m_{y}\neq 0$ in the large $r$ limit. One can further calculate the derivatives of the order parameters $O_{x}$ and $m_{y}$ with respect to the parameter $\lambda$ to determine the phase boundaries. Typical numerical results of $\partial O_{x}/\partial\lambda$ and $\partial m_{y}/\partial\lambda$ for $r=1000$ (which is large enough to approach the thermodynamic limit) for the Hermitian and non-Hermitian cases are shown in Figs. 5(h) and 5(i), respectively. The non-analytic points of the derivatives with respect to $\lambda$ correspond to the critical points. In this way, we determine the critical points from the order parameters in the whole $\lambda$-$\Gamma$ plane and plot them (blue points) in the phase diagram in Fig. 5(g).
To show that the critical points in the non-Hermitian cluster Ising model can be revealed by the quantum metric, we calculate $g_{\lambda\lambda}$ of the ground states as a function of $\lambda$ in Figs. 5(h) and 5(i). The locations of the peaks of $g_{\lambda\lambda}$ correspond to the critical points. We also plot these critical points (red points) in the phase diagram in Fig. 5(g). One can find that the critical points revealed via the quantum metric are consistent of those determined by the order parameters. However, the quantum metric is unable to witness the transition point between the quasi-long-ranged cluster (region III) and antiferromagnetic (region IV) phases, both of which have an imaginary gap. This invalidity is due to the absence of gap closing at the discontinuous transition points between phase regimes III and IV, similar to the case for the Hermitian p-wave-pair Aubry-André-Harper model Lv et al. (2022b). Notably, the critical points with gap closing can also be detected by the quantum metric $g_{\Gamma\Gamma}$. For instance, we plot $g_{\Gamma\Gamma}$ as a function of $\Gamma$ in Fig. 5(j). One can still find the peaks at $\Gamma\approx 2.23$ with fixed $\lambda=0.5$, and at $\Gamma\approx 0.48$ and $\Gamma\approx 2.95$ with fixed $\lambda=0.9$ in $g_{\Gamma\Gamma}$, indicating one and two critical points, respectively.
@@
Figure 6: (Color online) The non-Hermitian mixed-field Ising model $H_{\text{NH-Ising}}^{(2)}$ with the system size $N=10$. (a) $|M_{z}|$ and (b) $\xi_{h_{z}}$ of the ground state as functions of $h_{z}$ and $h_{x}$ under the PBC. The phase diagram based on $|M_{z}|$ consists two regions $\mathrm{I}$ and $\mathrm{II}$ for the paramagnetic (PM) and ferromagnetic (FM) phases with real and complex energies, respectively. $|M_{z}|$ and $\xi_{h_{z}}$ as a function of $h_{z}$ (c) under the PBC; and (d) the OBC, respectively. The auxiliary lines at $h_{z}/J=0.9$ (c) and $h_{z}/J=0.96$ (d) are plotted, respectively. Other parameter is $J=1$.